OpenStudy (anonymous):

Solve, using factoring: 8x2 = 16x

OpenStudy (campbell_st):

it is 8x(x - 2) = 0 the x = 0 or 2

OpenStudy (anonymous):

Got it thanks

OpenStudy (zepp):

You can simplify that before factoring \(8x^2\div 8=16x\div8\) \(x^2=2x\) Then \(x^2-2x=0\) \(x(x-2)\)=0